OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 4x + y = -1 1)(3, 1) 2)(-1, 3) 3)(-1, -3) 4)(1, 3)

OpenStudy (squirrels):

Are you familiar with the elimination method?

OpenStudy (anonymous):

yeah im just not very good at it

OpenStudy (squirrels):

Ok first thing you want to do is multiply -1 to the bottom equation so you can "eliminate" the bottom variable. \[-1(-2x+y=1)\] \[2y-y=1\]

OpenStudy (squirrels):

Oops that was the top equation, but it doesn't really matter which one you do Now you can cancel out the y variables \[2x-y=-1\]\[4x+y=-1\] \[6x=2\]

OpenStudy (anonymous):

simply subtract both of these..you will get -6x=2 =>x=-1/3 now reput it in any of given eq you will get.. say eq1 =>>-2(-1/3)+y=1 =>2/3+y=1 =>y=1-2/3 =>y=(3-2)/3 =>y=1/3 you will get. (x,y)=(-1/3,1/3) so your eq have some mistake,,,if answer is from those choices..!!!

OpenStudy (princeharryyy):

multiply equation 1 by number 2 which gives equation as 2(-2x + y = 1) =>-4x + 2y =2 add this equation to the second one 4x + y = -1 you get 3y=1 => y=1/3 and by replacing y in any equation u get x=-1/3

OpenStudy (princeharryyy):

the options are wrong check again

OpenStudy (anonymous):

@princeharryyy i have already mention that eq have some deviation from orignal values..!!1now it has been clear that eq are in -correct..

OpenStudy (squirrels):

ugh...

OpenStudy (princeharryyy):

@zaibali.qasmi question isn't wrong only the answers options are wrong )) hehe !!! enjoy !!! brother

OpenStudy (anonymous):

their right i copied and pasted them right off the parctice sheet

OpenStudy (princeharryyy):

may be they have given the wrong options in your practice sheet.... it's okay the answer is(-1/3,1/3) its 100% correct sweetie....

OpenStudy (anonymous):

sweetie... );;

OpenStudy (anonymous):

huh?